OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. If P = 104, and l = 13, what is the value of w?

OpenStudy (anonymous):

Put the values of P and l in the equation.... and solve. 104= 26 + 2w Now can you solve?

OpenStudy (anonymous):

104 = 26 + 2w 2w = 78 Divide by 2 both the sides and find w from here..